OpenStudy (wilder.monday):

Expand the following using the binomial theorem and Pascal’s triangle. 4.(2x – 3y)4 5.In the expansion of (3a + 4b)8, which of the following are possible variable terms? Explain your reasoning. a^2b^3; a^5b^3; ab^8; b^8; a^4b^4; a^8; ab^7; a^6b^5

OpenStudy (wilder.monday):

@ganeshie8 I think if I do 4 I'll be able to do 5.

ganeshie8 (ganeshie8):

for #4, use the fourth row http://ptri1.tripod.com/ptreal1r.gif

ganeshie8 (ganeshie8):

\[\begin{align} (2x-3y)^4 &= \bbox[3px, border:2px solid black]{?} (2x)^4(-3y)^0 + \bbox[3px, border:2px solid black]{?} (2x)^3(-3y)^1 + \bbox[3px, border:2px solid black]{?} (2x)^2(-3y)^2 \\~\\&+ \bbox[3px, border:2px solid black]{?} (2x)^1(-3y)^3+ \bbox[3px, border:2px solid black]{?} (2x)^0(-3y)^4 \end{align}\]

OpenStudy (wilder.monday):

okay so did it but then I was going to do what we did with the last question but I don't know what to do with the y ?

ganeshie8 (ganeshie8):

fill in the little squares with the coefficients from 4th row of pascal triangle

ganeshie8 (ganeshie8):

leave y as it is

ganeshie8 (ganeshie8):

it is just another variable like x

OpenStudy (wilder.monday):

so the first one would be -48yx^4

ganeshie8 (ganeshie8):

careful, (-3y)^0 = 1

ganeshie8 (ganeshie8):

anything^0 = 1

OpenStudy (wilder.monday):

so then it would just be 16x^4 ?

ganeshie8 (ganeshie8):

\[\begin{align} (2x-3y)^4 &= \bbox[3px, border:2px solid black]{1} (2x)^4(-3y)^0 + \bbox[3px, border:2px solid black]{4} (2x)^3(-3y)^1 + \bbox[3px, border:2px solid black]{6} (2x)^2(-3y)^2 \\~\\&+ \bbox[3px, border:2px solid black]{4} (2x)^1(-3y)^3+ \bbox[3px, border:2px solid black]{1} (2x)^0(-3y)^4 \end{align}\]

ganeshie8 (ganeshie8):

Yes! first term simplifies to 16x^4 good, keep going..

OpenStudy (wilder.monday):

the second one would be -24yx^3 ?

ganeshie8 (ganeshie8):

\[\large 4(2x)^3(-3y)^1\] this one ?

OpenStudy (wilder.monday):

yes, is what I got wrong?

ganeshie8 (ganeshie8):

lets see \[\large 4(2x)^3(-3y)^1\] first work the parenthesis \[\large -4(2^3x^3)(3^1y^1)\] \[\large -4(8x^3)(3y)\] \[\large -96x^3y\]

OpenStudy (wilder.monday):

so then it would be -216x^2y^2 ?

ganeshie8 (ganeshie8):

it would be +216x^2y^2 because (-1)^2 = 1

ganeshie8 (ganeshie8):

try remaining terms :)

OpenStudy (wilder.monday):

I don't think what I got for this one is right because I got -216xy^3

ganeshie8 (ganeshie8):

thats right!

ganeshie8 (ganeshie8):

finish off the last term

OpenStudy (wilder.monday):

oh okay :) the last one is 81y^4

ganeshie8 (ganeshie8):

You got it!

OpenStudy (wilder.monday):

yay thank you so much!!

ganeshie8 (ganeshie8):

yw :)

ganeshie8 (ganeshie8):

for #5, just pick the terms whose degree adds up to 8

OpenStudy (wilder.monday):

thanks :)

ganeshie8 (ganeshie8):

for example ` a^2b^3` cannot be in the expansion because the degree here is 2+3 = 5 which is not equal to 8.